OpenStudy (anonymous):

Which equation represents the line that passes through (2,4) and is perpendicular to the graph of y = 1/3 x -1

OpenStudy (anonymous):

the required line is (y-4)=-3(x-2) or y-4=-3x+6 or 3x+y=10

OpenStudy (anonymous):

on the last one is the 10 a positive?

OpenStudy (anonymous):

oh and here are the answer choices a. y = 1/3 = x + 10 b. y = 3x + 10 c. y = -3x - 2 d. y = -3x + 10

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thank you :)